OpenStudy (anonymous):

Which is not a solution to the equation x-2y=10 ? A. (–6, –8) B. (–3, 4) C. (8, –1) D. (14, 2)

OpenStudy (anonymous):

@YoungStudier

OpenStudy (youngstudier):

Ok, let's do what we did with the last one. Plug each answer choice into the equation :)

OpenStudy (anonymous):

ok ill start w/ A. (-6)-2(-8)=10. -6-2 is -8 *-8 is 64 so nope

OpenStudy (youngstudier):

Wait, hold up. Follow the order of operations. x - 2y = 10 -6 - 2(-8) = 10 Multiply first Then subtract, ok? :)

OpenStudy (anonymous):

ok..sorry for long time w/ out reply

OpenStudy (youngstudier):

It's alright. :)

OpenStudy (anonymous):

im stuck :/

OpenStudy (youngstudier):

Ok :) So the equation currently is: -6 - 2(-8) = 10 First, we wanna multiply 2 and -8 Next, we wanna subtract whatever 2 and -8 is from -6 If the ending result of the equation is 10, it's not the answer since we're looking for the choice that is NOT a solution to the equation.

OpenStudy (anonymous):

oh ok so 2*-8 is -16...

OpenStudy (anonymous):

and -6 - -16 is -22

OpenStudy (anonymous):

right?

OpenStudy (youngstudier):

o.o Nope -6 - (-16) is the same as -6 + 16 That would equal: ?

OpenStudy (anonymous):

-10 oh lol my bad

OpenStudy (youngstudier):

Positive 10 ;P Anyways, plug B in next. :)

OpenStudy (anonymous):

10 dang im bad with math

OpenStudy (youngstudier):

Ok, so let's plug B in next: x - 2y = 10 -3 - 2*4 = 10 ? = 10

OpenStudy (anonymous):

-3-2(4)=-11

OpenStudy (youngstudier):

Correct. Do you think that's our answer?

OpenStudy (anonymous):

ya

OpenStudy (youngstudier):

Yeppers. ^.^ So B is our answer. :)

OpenStudy (anonymous):

k thanks!:)

OpenStudy (youngstudier):

No problem! ^_^